1. For the following cashflows, compute the IRR: Year Cashflows -$14,000 5,000 5,500 6,000 6,500PLEASE WORK OUT DO NOT USE EXCEL
Answer #1

let i% be the IRR, then

PW = -14000 + 5000 * (P/A,i%,4) + 500 * (P/G,i%,4) = 0

5000 * (P/A,i%,4) + 500 * (P/G,i%,4) = 14000

dividing by 500, we get

10 * (P/A,i%,4) + (P/G,i%,4) = 28

using trail and error method

When I = 20%, value of 10 * (P/A,i%,4) + (P/G,i%,4) = 10 *2.588735 + 3.298611 = 29.1859

When I = 21%, value of 10 * (P/A,i%,4) + (P/G,i%,4) = 10 *2.540441 + 3.211483 = 28.6159

When I = 22%, value of 10 * (P/A,i%,4) + (P/G,i%,4) = 10 *2.493641 + 3.127473 = 28.0639

When I = 23%, value of 10 * (P/A,i%,4) + (P/G,i%,4) = 10 *2.448272 + 3.046443 = 27.5292

using interpolation

i = 22% + (28.0639 - 28) / (28.0639 - 27.5292)*(23% - 22%)

i = 22% + 0.1195%

i = 22.12% (Approx)